OpenStudy (anonymous):

ind the equation of a line given the point and slope below. (6, 6) no slope

OpenStudy (anonymous):

if m=0 (not sure if that's what you meant by no slope), then y=mx+b plug in your 6 for y and 6 for x 6=(0)(6)+b b=6 y=6 is your equation

OpenStudy (jdoe0001):

no slope could also mean, undefined?